### v1

# When Video Coding Meets Multimodal Large Language Models: A Unified Paradigm for Video Coding

Abstract
Existing codecs are designed to eliminate intrinsic redundancies to create a compact representation for compression. However, strong external priors from Multimodal Large Language Models (MLLMs) have not been explicitly explored in video compression. Herein, we introduce a unified paradigm for Cross-Modality Video Coding (CMVC), which is a pioneering approach to explore multimodality representation and video generative models in video coding. Specifically, on the encoder side, we disentangle a video into spatial content and motion components, which are subsequently transformed into distinct modalities to achieve very compact representation by leveraging MLLMs. During decoding, previously encoded components and video generation models are leveraged to create multiple encoding-decoding modes that optimize video reconstruction quality for specific decoding requirements, including Text-Text-to-Video (TT2V) mode to ensure high-quality semantic information and Image-Text-to-Video (IT2V) mode to achieve superb perceptual consistency. In addition, we propose an efficient frame interpolation model for IT2V mode via Low-Rank Adaption (LoRA) tuning to guarantee perceptual quality, which allows the generated motion cues to behave smoothly. Experiments on benchmarks indicate that TT2V achieves effective semantic reconstruction, while IT2V exhibits competitive perceptual consistency. These results highlight potential directions for future research in video coding.
Introduction
The objective of video compression is to eliminate redundancies, thereby reducing the overall volume of video data at the same quality level. In recent years, traditional and deep video codecs have primarily focused on preserving signal-level quality (Li, Li, and Lu 2021, 2023). However, existing codecs often fall short of preserving essential semantic information at high compression ratios, leading to rather low-quality video reconstruction. Fortunately, the rapid advancement of Multimodal Large Language Models (MLLMs) has been opening up promising possibilities for Ultra- and Extremely- Low Bitrate (ULB and ELB) compression.
MLLMs are advanced AI models designed to process and understand data from multiple modalities, such as text, images, and audio. Many image compression work, such as VR-CMC (Li et al. 2021), SCMC (Zhang et al. 2023), MCM (Gao et al. 2024), and CMC-Bench (Li et al. 2024), etc., leverage the capabilities of cross-modality representation. These Cross-Modality Image Compression (CMIC) techniques first transmute images into textual (I2T) formats and then produce dynamic content with similar semantics. This transformation not only distills essential semantic content but also significantly reduces the storage requirements through compression compared to the original image data. However, the integration of cross-modality representation and MLLMs in the domain of video coding is still relatively underexplored.
Figure 1: The framework of the proposed CMVC scheme. We separate videos into different video clips through the keyframe selection strategy and extract the content of keyframe and motion from video clips. Then, MLLMs are employed to create multimodal representations for these components. These representations of keyframes and motion are then compressed and transmitted to the decoder. We provide two exemplary modes, including TT2V and IT2V modes, for video generation. This model offers flexibility to integrate various SOTA models and mode conversions while maintaining semantic and perceptual quality at high compression ratios.
Video coding, compared with image compression, additionally requires an exploration of how to effectively represent and compress temporal information. Fortunately, MLLMs (Lin et al. 2023; Zhang, Li, and Bing 2023; Fei et al. 2024) possess extensive knowledge relevant to video content analysis, including an understanding of context, objects, and actions within the video. As a result, MLLMs excel at handling sequential data and understanding the temporal relationships of events in videos. Equipped with these capabilities, MLLMs are particularly well-suited for the effective extraction of temporal information for video tasks. In this manner, MLLMs can be utilized as powerful tools to extract compact textual representations from videos based on specific prompts.
Through integrating cutting-edge MLLMs, we propose a new Cross-Modality Video Coding (CMVC) paradigm to optimize the representation of both content and motion with video generation models for video coding. Depending on the intriguing nature of multimodality representations, various encoding-decoding modes can be further established. We propose two exemplary modes alternatively: TT2V (Text-Text-to-Video) mode for semantic reconstruction at ULB and IT2V (Image-Text-to-Video) mode for perceptual reconstruction at ELB. In the TT2V mode, inspired by the workflow of CMIC and the stunning generation capability of existing TT2V models for video reconstruction, we first extract the representative text constructed with our selection strategy, effectively encoding video content and motion. Then, the video generation model is utilized to reconstruct the corresponding video from text inputs. The rationale behind this strategy lies in that the compact yet effective text representations from the encoder encapsulate semantic details to enable high-quality semantic reconstruction for the decoder. Different from TT2V, the IT2V mode is designed to enhance perceptual reconstruction, since images provide richer visual contexts compared to text, benefiting perceptual consistency. This is achieved by inputting similar text representations with TT2V mode and extra selected keyframes from the encoder to the decoder for better perceptual video reconstruction. To further ensure perceptual smoothness across consecutive frames, an efficient adaptation tuning in a frame-interpolation manner via Low-Rank Adaption (LoRA) tuning is tailored to fully exploit the semantic cues and visual contexts from both input texts and keyframes to facilitate high-quality perceptual consistency for video reconstruction. This comprehensive paradigm adeptly accommodates diverse modality representations within video coding, by tapping into foundational MLLMs and video generation models, which sheds light on future video coding works. The contributions of our work are as follows:
  • We introduce a unified paradigm for CMVC. To the best of our knowledge, this is the first work to make utility of foundational MLLMs and video generation models for video coding.
  • We elaborate multiple encoding-decoding modes to achieve good trade-off video reconstruction quality depending on specified decoding requirements, including TT2V mode to ensure high-quality semantic information and IT2V mode to achieve superb perceptual consistency.
  • Extensive experiments demonstrate that our proposed CMVC pipeline obtains competitive video reconstructions on HEVC Class B, C, D, E, UVG and MCL-JCV benchmarks, while also maintaining high compression rates.
Related works
Video Generation Models
Recently, video generation models have emerged as an increasingly promising topic, with numerous studies (Ho et al. 2022; Singer et al. 2023; Ge et al. 2023; Blattmann et al. 2023; Chen et al. 2024a; Wang et al. 2023a) showcasing promising advancements, that enables generative models simulate the real world principle. These include various approaches such as text-to-video (T2V) (Wang et al. 2023a, b), image-to-video (I2V) (Chen et al. 2024b; Esser et al. 2023; Yin et al. 2023), and image-text-to-video (IT2V) (Zhang et al. 2024a), among others. T2V technology is designed to convert descriptive text into corresponding videos (Lin et al. 2023; Wang et al. 2023a; Zhang, Li, and Bing 2023). One of the primary challenges in this field is to understand the intricate semantics of the input text and effectively translate it into dynamic visual content, following real-world physics. To achieve optimal quality in the generated videos, these models are trained on large-scale video datasets, leveraging a large text-video corpus to train the model for better alignment. I2V generative models typically include methods such as video interpolation and image-driven video diffusion. Image-driven video diffusion models (Voleti et al. 2024; Chai et al. 2023; Ouyang et al. 2024) necessitate the given referring image to steer the generative model to produce corresponding videos. Compared to image-driven video diffusion models, video interpolation techniques (Huang et al. 2022; Li et al. 2023) can better maintain consistency in both resolution and motions in terms of moving objects across consecutive frames, which elucidates great potential utility for image-driven video generation. In contrast to I2V models, IT2V generative models incorporate textual guidance to enhance video generation. For instance, DiffMorpher (Zhang et al. 2024a) adds visual descriptions for images by adopting latent interpolation adaptation training to produce smoothing transformation. Building on this concept, we can also enhance video generation by incorporating temporal descriptions.
Cross-Modality Compression
Multimodal generation has been effectively applied in the field of compression (Lu et al. 2022). To preserve semantic communication at ELB, CMC (Li et al. 2021) integrates the I2T translation model with the T2I (Text-to-Image) generation model. Building on this foundation, SCMC (Zhang et al. 2023) introduces a scalable cross-modality compression paradigm that hierarchically represents images across different modalities, thereby enhancing both semantic and signal-level fidelity. Subsequently, VR-CMC (Gao et al. 2024), a variable-rate cross-modal compression technique, employs variable-rate prompts to capture data at varying levels of granularity. Additionally, a CMC benchmark has been established for image compression (Li et al. 2024). These models demonstrate that the integration of I2T and T2I methodologies has outperformed the most advanced visual signal codecs. Despite these advancements, there remains limited research focused on cross-modality video coding
Figure 2: Visualization of the results from the keyframe selection strategy. The red points indicate the index of the keyframes. The first and last frames are set as keyframes by default.
CMVC Scheme
Overview
We propose a unified paradigm for cross-modality video compression (CMVC), as illustrated in Fig. 1. First, the video is segmented into multiple clips through the keyframe selection strategy, and each clip is disentangled into the contents of keyframes and motion components. Subsequently, MLLMs are employed to create multimodal representations for these components. For instance, keyframes and motion can be represented using V2T models to convert them into textual formats. These representations of keyframes and motion are then compressed and transmitted to the decoder. We design two distinct modes based on the modalities of keyframes and motion: TT2V mode and IT2V mode. In TT2V mode, keyframes and motion are both represented as text, enabling the generation of videos with high semantic consistency at ULB. In IT2V mode, keyframes are represented as images, and motion is described using text, facilitating video generation with high perceptual consistency. Overall, this paradigm offers flexibility to integrate various SOTA models and mode conversions while maintaining semantic and perceptual quality at high compression ratios.
CMVC Encoder
Keyframe selection strategy. Keyframes divide a full length video sequence into clips. Let $n$ denote the number of keyframes, allowing us to extract $n$-1 clips from the video, with the first and last frames initially designated as keyframes. The first frame is encoded using the CLIP encoder (Radford et al. 2021) to extract a high-level feature vector $v_{k}$ containing concise semantic information. We calculate the cosine similarity distance between the first frame and subsequent frames as follows:
$\centering\mathcal{D}_{v}=\frac{v_{k}\cdot v_{k+i}}{\left\|v_{k}\right\|\cdot\left\|v_{k+i}\right\|}=\frac{\sum_{j=1}^{m}v_{k,j}\cdot v_{k+i,j}}{\sqrt{\sum_{j=1}^{m}\left(v_{k,j}\right)^{2}\cdot\sum_{j=1}^{m}\left(v_{k+i,j}\right)^{2}}},\@add@centering$ (1)
where $v_{k+i}$ is the feature vectors extracted from the subsequent frames. $m$ is the number of components of vectors $v_{k}$ and $v_{k+i}$. Within a uniform interval, we select the frame with the smallest similarity to the previous keyframes to form a set of keyframes that better showcase significant motion. Then, the $v_{k}$ is replaced by the next keyframe features, acting like a dynamic mechanism. This iterative process is repeated for subsequent clips, systematically identifying representative keyframes. We display the result of the keyframe selection strategy in Fig. 2.
Multimodality representation. In our proposed scheme, we focus on efficiently representing spatial and temporal information of videos through keyframes and motion. Specifically, let $V$ represent the original video. The keyframes are denoted as $K$ = {$k_{1}$, $k_{2}$, …, $k_{n}$}, where each $k_{i}$ is a keyframe. The motion information $m_{j}$ between these keyframes is represented by $M$ = {$m_{1}$, $m_{2}$, …, $m_{n-1}$}. Keyframes and motion are transformed into multimodality representations as follows: $T_{k,i}=f(k_{i})$ and $T_{m,j}=g(m_{j})$. Here, $f(*)$ and $g(*)$ denote the process of cross-modality representation for keyframes and motion, respectively. As illustrated in Fig. 1, keyframes can be transformed into textual and visual image representations, while motion can be represented by textual description and optical flow. Thus, the total bitrate is given by:
$R_{total}=\sum_{i=1}^{n}R_{k}(T_{k,i})+\sum_{j=1}^{n-1}R_{m}(T_{m,j}),$ (2)
where $R_{k}$ and $R_{m}$ are the entropy coding modules for keyframes and motion, respectively. The bitrates can be adjusted by $n$ and the compression ratio of keyframes and motion.
CMVC Decoder
In the decoder, we utilize decoded keyframe $\hat{K}$ and motion $\hat{M}$ to achieve video generation, as follows:
$\hat{V}=\mathcal{G}(\hat{K},\hat{M}),$ (3)
where $\mathcal{G}(*)$ is a video generation model and $\hat{V}$ is the reconstructed video. According to different modality representations for keyframe and motion, we design two modes, including TT2V mode and IT2V mode.
In the TT2V mode, we employ SOTA video generation models, such as Open-Sora (Lab and etc. 2024), VideoCrafter2 (Chen et al. 2024a), AnimateDiff (Guo et al. 2024) and etc., to generate videos based on the decoded descriptions of keyframes and motion. By leveraging advancements in these models, we can employ a range of techniques that lead to optimal semantic reconstruction. Our findings indicate that more detailed descriptions correspond to higher bitrates, resulting in better semantic quality of the generated videos. In the IT2V mode, we integrate keyframe images and motion descriptions to enhance the perceptual quality of the generative video. In addition to adopting existing IT2V models, we propose an efficient generative model utilizing LoRA tuning to ensure higher-quality perceptual consistency at ELB.
The IT2V generative model.
Figure 3: The workflow of the IT2V generative model. Two LoRAs are trained to fit the two keyframe images ($I_{0}$ and $I_{1}$), respectively. To generate $w$-th frame between $I_{0}$ and $I_{1}$, we interpolate $I^{{}^{\prime}}_{w}$ and the LoRA parameters according to the weights $w_{i}$ and $w_{l}$, where $w_{i}$ and $w_{l}$ are optimized according to the $w$-th frame.
The IT2V mode is designed to obtain a reconstructed video according to keyframe images and the text of motion. Thus, we propose a IT2V generative model, which generates a video clip according to two keyframe images ($I_{0}$ and $I_{1}$) and the description of the motion of this video clip. Specifically, we adopt a stable diffusion (SD) model with LoRA, which fine-tunes the model parameters $\theta$ by training a low-rank residual component $\triangle\theta$. This residual can be decomposed into products of low-rank matrices. LoRA demonstrates significant efficiency in generating various samples while maintaining consistent semantic identity across different latent noise traversals. The proposed IT2V model is shown in Fig. 3. We first train two LoRAs ($\Delta\theta_{0}$ and $\Delta\theta_{1}$) on the SD UNet $\epsilon_{\theta}$ for each of the two images $I_{0}$ and $I_{1}$. The learning objective of $\Delta\theta_{i}(i=0,1)$ is:
$\mathcal{L}\left(\Delta\theta_{i}\right)=\mathbb{E}_{\epsilon,t}\left[\left\|\epsilon-\epsilon_{\theta+\Delta\theta_{i}}\left(\mathbf{z}_{t,i},t,\mathbf{c}_{i}\right)\right\|^{2}\right],$ (4)
where $\mathbf{z}_{t,i}=\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{z}}_{i}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ is the noised latent embedding at diffusion step $t$. $\hat{\mathbf{z}}_{i}$ is the VAE encoded latent of the $I_{i}$ image. $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the random sampled Gaussian noise. $\mathbf{c}_{i}$ is the motion embedding encoded from the motion prompt. $\epsilon_{\theta+\Delta\theta_{i}}$ represents the LoRA-integrated UNet. The fine-tuning objective is optimized separately via gradient descent in $\Delta\theta_{0}$ and $\Delta\theta_{1}$.
Figure 4: Left: Comparison results of combination of different V2T models (VideoLLaVA and VideoLLaMA) and TT2V models (VideoCrafter1, VideoCrafter2, ModelScope, OpenSora and AnimateDiff). Right: Visual quality comparison of the TT2V mode and VVC. At ULB, our proposed TT2V mode successfully preserves the semantic quality of the videos. In contrast, VVC brings significant blocking artifacts, which impedes the effective conveyance of semantic information in videos.
Figure 5: Visual quality comparison. The values represent the BPP (1e-2) and the DISTS value. A lower DISTS value indicates better perceptual quality.
Frame and model interpolation. In order to generate $I^{{}^{\prime}}_{w}$, we first conduct keyframe interpolation as input in the following manner:
$I_{w}^{\prime}=w_{i}\times I_{0}+\left(1-w_{i}\right)\times I_{1}.$ (5)
Building upon DiffMorpher (Zhang et al. 2024a), we then interpolate the model weight $\Delta\theta_{l}$ according to $\Delta\theta_{0}$ and $\Delta\theta_{1}$:
$\Delta\theta_{l}=w_{l}\times\Delta\theta_{0}+\left(1-w_{l}\right)\times\Delta\theta_{1}.$ (6)
$\Delta\theta_{l}$ is the LoRA parameters, which are integrated to UNet $\epsilon_{\theta+\Delta\theta_{l}}$.
The uniformly linear interpolation schedule may result in an uneven transition. Thus, we conduct online training for the $w_{i}$ and $w_{l}$ with the constraint of $D(I_{w},\hat{I}_{w})$ at the encoder side, where $D(*)$ is the $\mathcal{L}_{2}$ loss. We only update the $w_{i}$ and $w_{l}$, as follows:
$w^{t+1}_{i}=w^{t}_{i}-\alpha\nabla D(w^{t}_{i}),$ (7)
$w^{t+1}_{l}=w^{t}_{l}-\alpha\nabla D(w^{t}_{l}),$ (8)
where $w^{t}_{i}$ and $w^{t}_{l}$ is the parameter $w_{i}$ and $w_{l}$ at training step $t$, resepctively. $\alpha$ is the learning rate, set to 0.001, while $\nabla D(*)$ denotes the gradient of the loss function concerning the parameters at the training step $t$. After obtaining optimal $w^{t}_{i}$ and $w^{t}_{l}$, we compress and transmit them to the decoder. The VAE decoder then reconstructs the denoised latent representation into the $w$-th frame, resulting in $\hat{I}_{w}$.
Experiments
Experimental settings
Datasets. The datasets, including HEVC Class B, C, D, and E, as well as UVG and MCL-JCV, are extensively utilized for evaluating both traditional and neural video codecs. These datasets vary in resolution and content, providing a diverse range of scenarios for comprehensive assessment. To ensure compatibility with various video codecs, we resize videos to dimensions that are multiples of 64 for both width and height.
Figure 6: The R-D performance comparison in the IT2V mode. The comparisons are performed on the Class B, Class C, Class D, Class E, UVG, and MCL-JCV, respectively.
Table 1: BD-Rate (%) comparison of different video generation models across various datasets in terms of DISTS.
RIFE AMT DiffMorpher Ours
Class B 6.00 47.90 -46.47 -65.71
Class C 21.85 14.87 -$3$31.15
Class D -17.82 -29.88 -51.40 -54.45
Class E -24.61 -79.22 38.03 -90.40
UVG 23.87 42.59 13.62 -24.21
MCL-JCV 87.37 164.10 24.60 -19.73
Comparison methods. There are numerous SOTA foundation models available for video understanding. We choose two prominent models, namely VideoLLaVA and VideoLLaMA, to extract semantic information from videos. This process aligns with the V2T stage depicted in Fig. 1, where the selected models play an important role in extracting semantic descriptions for keyframes and motion. In the TT2V mode, numerous video generation models are available. In this context, we employ advanced video generation models, including Open-Sora, VideoCrafter1, VideoCrafter2, and AnimateDiff, for the purpose of generating videos based on textual input. In addition, we compare with the video codec VVC at the extremely low bitrate with QP=63. In the IT2V mode, we conduct a comparative analysis of existing traditional video codecs, such as x264, x265, and VVC (Bross et al. 2021). Alongside this, we evaluate our method against deep video codecs such as DCVC (Li, Li, and Lu 2021) and DCVC-DC (Li, Li, and Lu 2023), but these codecs encounter challenges in achieving extremely low bitrate coding. In addition, we compare the video generation technique, DiffMorpher (Zhang et al. 2024b), which requires keyframe images and motion descriptions for controlling video generation. In our exploration of various video interpolation methods, it is essential to note that these approaches rely solely on keyframes for control, omitting any incorporation of motion descriptions. Furthermore, it should be emphasized that the bit consumption associated with motion text has not been calculated. Notably, the anchor of BD-Rate is VVC under Low Delay P configuration with QPs={63, 52, 47, 42}.
Figure 7: Visual quality comparison in a video. The numbers displayed beneath the images correspond to the frame index.
Figure 8: Visual quality comparison between the TT2V mode and the IT2V mode. The TT2V mode effectively preserves semantic consistency with the ground truth, while the IT2V mode is designed to keep perceptual consistency.
Experimental results
Comparison in the TT2V mode. We conduct a comparative analysis of two V2T models, VideoLLaVA and VideoLLaMA, both of which are state-of-the-art (SOTA) MLLMs. Subsequently, we extend our comparison to include five video generation models: VideoCrafter1, VideoCrafter2, ModelScope, OpenSora, and AnimateDiff. Furthermore, we compare these models against the traditional video codec VVC at QP=63, which results in a higher bitrate than our proposed scheme. Our assessment focuses on five aspects: subject consistency, background consistency, temporal flickering, motion smoothness, and frame quality. The results, illustrated in Fig. 4, indicate that the TT2V generation models outperformed VVC, showcasing better frame-wise quality and consistency in both background and subject representation. These results reflect the average performance across all testing datasets, and detailed comparison results can be found in the supplementary material. The visual quality comparisons illustrated in Fig. 4 indicate that VVC displays considerable blocking artifacts, which severely hinder its ability to convey semantic information.
Table 2: Ablation studies on the keyframe selection strategy.
Models Settings BD-Rate(%) $\downarrow$
Sampling strategies Uniform sampling -19.705
Random sampling -9.859
MSE -14.496
CS -24.206
Keyframe number 2 -20.001
3 -5.104
4 12.665
Keyframe quality low -24.206
middle -12.043
high -2.617
Comparison in the IT2V mode. We compare our model with traditional codecs (x264, x265, and VVC) as well as deep video codecs (DCVC and DCVC-DC). As presented in Fig. 6, we use DISTS to evaluate perceptual quality. Additional comparisons with other evaluation metrics, such as LPIPS, FID, and PSNR, are provided in the supplementary material. However, the pretrained models provided by deep video codecs have limitations in achieving ELB. In addition, we compare our proposed model with various video generation models, including RIFE (Huang et al. 2022), AMT (Li et al. 2023), and DiffMorpher (Zhang et al. 2024a), as detailed in Table 1. By adjusting the number and quality of keyframe images, we can effectively control the bitrate. For our comparisons, we select the optimal results for comparison, where the settings can be found in the supplementary material. Our model exhibits superior performance across most datasets, demonstrating greater stability compared to other video generation models. The visual quality is evaluated at similar bitrates, as shown in Fig. 5 and Fig. 7. Our proposed model exhibits superior perceptual quality in both spatial and temporal dimensions. Additionally, we showcase frames sampled from the decoded videos generated by the TT2V mode and the IT2V mode. The TT2V mode effectively preserves semantic consistency with the ground truth, while the IT2V mode further ensures perceptual consistency.
Ablation studies
Keyframe. We perform ablation studies focused on keyframes, examining various aspects such as keyframe selection methods, the quality of keyframe images, and the number of keyframe images. In our keyframe selection process, we evaluate various sampling strategies, including uniform sampling and random sampling. Given that these techniques do not rely on a distance function, we also compare the sampling strategy with mean-square error (MSE) distance to Ours with cosine similarity (CS) distance. Regarding the quality of keyframe images, we varied the quality levels, including low, medium, and high quality, which correspond to the compression factors 64, 128, and 256, respectively. The results presented in Table 2 indicate that higher quality decoded images result in increased bitrate consumption, such that higher quality does not necessarily lead to a better BD-Rate. Adjusting the number of keyframes based on the frame number of the video, we observe that a lower number of keyframes can maintain a balance between quality and bitrate consumption.
IT2V generative model. We perform ablation studies focused on the different settings, including the influence of motion description, different codecs, updating strategies, training step, and sampling step. In terms of motion description, we compare the model without motion description, as depicted in Table 3. The results indicate that incorporating motion description significantly enhances video reconstruction quality. Additionally, we explore a range of codecs for keyframe images, such as Hyperprior (Ballé et al. 2018), NIC (Chen et al. 2021), and NTVP. Among these, NTVP stands out by demonstrating superior reconstructed quality while maintaining a lower coding rate. Our model requires updating $w_{i}$ and $w_{l}$ based on the input, such that we further evaluate the effectiveness of updating strategies, as illustrated in Table 3. To assess the effectiveness of these updating strategies, we present further evaluations in Table 3. Moreover, we examine the repercussions of varying training and sampling steps. An increase in the number of sampling steps correlates with improved results. To strike a balance between performance and computational efficiency, we choose 100 training steps and 50 sampling steps for our final implementation.
Table 3: Ablation studies of the IT2V mode.
Models Settings BD-Rate(%) $\downarrow$
Motion description $\times$ 8.516
$\checkmark$ -24.206
Codecs Hyperprior -12.261
NIC -16.132
NTVP -24.206
Updating $w_{i}$$\times$ -12.402
Updating $w_{l}$$\times$ -18.551
Updating $w_{i}$ and $w_{l}$$\times$ -2.351
Updating $w_{i}$ and $w_{l}$checkmark -24.206
Training step 50 -17.864
100 -24.206
150 -25.196
Sampling step 20 -4.035
50 -24.206
100 -25.207
Conclusion and Discussion
We propose a CMVC paradigm that represents a promising advancement in video coding technology. This framework effectively tackles the challenges of preserving semantic integrity and perceptual consistency at ULB and ELB. By leveraging MLLMs and cross-modality representation techniques, the proposed CMVC framework disentangles videos into content and motion components, transforming them into different modalities for efficient compression and reconstruction. Through the TT2V and IT2V modes, CMVC achieves a balance between semantic information and perceptual quality, offering a comprehensive solution at high compression ratios.
CMVC includes the TT2V mode and IT2V mode, but more modes can be further explored. For instance, motion representation can be realized through optical flow or trajectories. By integrating multiple modalities of keyframes and motion, we can cater to diverse reconstruction requirements. Furthermore, future efforts should prioritize enhancing CMVC at higher bitrates by integrating more control information to facilitate the reconstruction of the original video. This approach aims to achieve superior performance across all bitrates and dimensions when compared to traditional codecs.

### v2

@@
 Abstract
-Existing codecs are designed to eliminate intrinsic redundancies to create a compact representation for compression.
-However, strong external priors from Multimodal Large Language Models (MLLMs) have not been explicitly explored in video compression.
-Herein, we introduce a unified paradigm for Cross-Modality Video Coding (CMVC), which is a pioneering approach to explore multimodality representation and video generative models in video coding.
-Specifically, on the encoder side, we disentangle a video into spatial content and motion components, which are subsequently transformed into distinct modalities to achieve very compact representation by leveraging MLLMs.
-During decoding, previously encoded components and video generation models are leveraged to create multiple encoding-decoding modes that optimize video reconstruction quality for specific decoding requirements, including Text-Text-to-Video (TT2V) mode to ensure high-quality semantic information and Image-Text-to-Video (IT2V) mode to achieve superb perceptual consistency.
-In addition, we propose an efficient frame interpolation model for IT2V mode via Low-Rank Adaption (LoRA) tuning to guarantee perceptual quality, which allows the generated motion cues to behave smoothly.
-Experiments on benchmarks indicate that TT2V achieves effective semantic reconstruction, while IT2V exhibits competitive perceptual consistency.
-These results highlight potential directions for future research in video coding.
+Traditional video compression methods perform well at high bitrates but struggle to preserve fine-grained semantic information at low bitrates.
+Recently, with the blossom of Multimodal Large Language Models (MLLMs), Cross-modal compression techniques offer prospective solutions for improving video compression under low-bitrate conditions.
+In this paper, we propose a unified Cross-Modality Video Coding (CMVC) framework that integrates multimodal representations and video generative models.
+The encoder disentangles video into spatial and temporal components, which are mapped to compact cross-modal representations using MLLMs.
+During decoding, different encoding-decoding modes are employed to acquire various video reconstruction quality, including Text-Text-to-Video (TT2V) for semantic preservation and Image-Text-to-Video (IT2V) for perceptual consistency.
+Additionally, we elaborate an efficient frame interpolation model using Low-Rank Adaptation (LoRA) to improve the perceptual quality.
+Experiments show that TT2V achieves effective semantic reconstruction, while IT2V ensures competitive perceptual consistency.
+These findings suggest the potential of leveraging multimodal priors to improve video compression, offering promising future research directions.
 Introduction
-The objective of video compression is to eliminate redundancies, thereby reducing the overall volume of video data at the same quality level.
-In recent years, traditional and deep video codecs have primarily focused on preserving signal-level quality (Li, Li, and Lu 2021, 2023).
-However, existing codecs often fall short of preserving essential semantic information at high compression ratios, leading to rather low-quality video reconstruction.
-Fortunately, the rapid advancement of Multimodal Large Language Models (MLLMs) has been opening up promising possibilities for Ultra- and Extremely- Low Bitrate (ULB and ELB) compression.
-MLLMs are advanced AI models designed to process and understand data from multiple modalities, such as text, images, and audio.
-Many image compression work, such as VR-CMC (Li et al.
+Traditional video compression (Bhaskaran and Konstantinides 1997; Bross et al.
+2021; Ma et al.
+2019) has primarily focused on signal-level reconstruction, optimizing components such as pixel intensities and motion vectors during encoding and decoding.
+This approach has yielded substantial advancements in high-bitrate scenarios, exemplified by codecs like VVC (Bross et al.
+2021) and DCVC (Li, Li, and Lu 2021), which deliver remarkable performance.
+However, these conventional methods often encounter limitations at low bitrates scenario, particularly when it comes to preserving fine-grained semantic information, crucial for applications such as disaster response and remote surveillance.
+In such contexts, reconstructed video is prone to the loss or blurring of key features.
+To address these shortcomings at low bitrates, cross-modal compression (CMC) methods have emerged (Li et al.
+2021; Zhang et al.
+2023; Gao et al.
+2024), which harness the complementary strengths of multiple modalities to enhance image representations under constrained bitrate conditions.
+Notable approaches, such as VR-CMC (Li et al.
 2021), SCMC (Zhang et al.
@@
 2024), and CMC-Bench (Li et al.
-2024), etc., leverage the capabilities of cross-modality representation.
-These Cross-Modality Image Compression (CMIC) techniques first transmute images into textual (I2T) formats and then produce dynamic content with similar semantics.
-This transformation not only distills essential semantic content but also significantly reduces the storage requirements through compression compared to the original image data.
-However, the integration of cross-modality representation and MLLMs in the domain of video coding is still relatively underexplored.
+2024), exploit the transformation of images into textual formats (I2T) to generate semantically similar content.
+This transformation not only preserves crucial semantic information but also achieves significant reductions in storage requirements, resulting in higher compression ratios compared to conventional image data representations.
+However, the potential of cross-modal representations for video compression, particularly in maintaining both semantic and perceptual quality at low bitrates, remains to be relatively underexplored.
+Unlike image compression, which interests primarily in spatial information, video compression additionally involves the representation of temporal dynamics, significantly increasing its complexity.
+In this regard, Multimodal Large Language Models (MLLMs) (Han et al.
+2022; Ruan et al.
+2023) offer distinct advantages due to their inherent capacity to analyze and interpret temporal relationships within video content.
+Recent researches (Lin et al.
+2023; Zhang, Li, and Bing 2023; Fei et al.
+2024) have demonstrated that MLLMs excel at processing sequential data and capturing dependencies across temporal events.
+By leveraging these capabilities, MLLMs can generate compact and semantically rich textual representations of video content, facilitating efficient compression while preserving high-quality reconstruction, even in low-bitrate settings.
+As such, MLLMs present a promising paradigm for advancing video compression, offering an effective balance between encoding efficiency and semantic preservation.
 Figure 1: The framework of the proposed CMVC scheme.
-We separate videos into different video clips through the keyframe selection strategy and extract the content of keyframe and motion from video clips.
-Then, MLLMs are employed to create multimodal representations for these components.
-These representations of keyframes and motion are then compressed and transmitted to the decoder.
-We provide two exemplary modes, including TT2V and IT2V modes, for video generation.
-This model offers flexibility to integrate various SOTA models and mode conversions while maintaining semantic and perceptual quality at high compression ratios.
-Video coding, compared with image compression, additionally requires an exploration of how to effectively represent and compress temporal information.
-Fortunately, MLLMs (Lin et al.
-2023; Zhang, Li, and Bing 2023; Fei et al.
-2024) possess extensive knowledge relevant to video content analysis, including an understanding of context, objects, and actions within the video.
-As a result, MLLMs excel at handling sequential data and understanding the temporal relationships of events in videos.
-Equipped with these capabilities, MLLMs are particularly well-suited for the effective extraction of temporal information for video tasks.
-In this manner, MLLMs can be utilized as powerful tools to extract compact textual representations from videos based on specific prompts.
-Through integrating cutting-edge MLLMs, we propose a new Cross-Modality Video Coding (CMVC) paradigm to optimize the representation of both content and motion with video generation models for video coding.
-Depending on the intriguing nature of multimodality representations, various encoding-decoding modes can be further established.
-We propose two exemplary modes alternatively: TT2V (Text-Text-to-Video) mode for semantic reconstruction at ULB and IT2V (Image-Text-to-Video) mode for perceptual reconstruction at ELB.
-In the TT2V mode, inspired by the workflow of CMIC and the stunning generation capability of existing TT2V models for video reconstruction, we first extract the representative text constructed with our selection strategy, effectively encoding video content and motion.
+This framework operates by first segmenting the video into distinct clips using a keyframe selection strategy (a), allowing for the extraction of both spatial and temporal components from each video segment.
+Subsequently, MLLMs are employed to generate multimodal representations of these components.
+For instance, spatial information can be represented through text or images, while temporal dynamics may be encoded using text or audio modalities.
+These multimodal representations are then encoded via their respective encoders, resulting in compressed bitstreams for each component.
+The bitstreams corresponding to different components are then combined and transmitted to the decoder.
+In the decoder, we provide two exemplary modes, including TT2V (c) and IT2V (d) modes, for video generation.
+This model integrate various SoTA models and mode conversions while maintaining semantic and perceptual quality at relatively high compression ratios.
+Motivated by the advantages of MLLMs, we propose a new Cross-Modality Video Coding (CMVC) scheme aimed at optimizing the representation of both spatial and temporal components, thereby enabling high-fidelity semantic and perceptual reconstruction at low bitrates.
+Capitalizing on the rich potential of multimodal representations, this framework supports the development of diverse encoding-decoding modes tailored to specific reconstruction requirements.
+We propose two exemplary modes alternatively: TT2V (Text-Text-to-Video) mode for semantic reconstruction at ultra-low bitrate (ULB) and IT2V (Image-Text-to-Video) mode for perceptual reconstruction at extremely low bitrate (ELB).
+In the TT2V mode, inspired by the workflow of Cross-Modality Image Compression (CMIC) (Li et al.
+2021; Zhang et al.
+2023; Li et al.
+2024) and the stunning generation capability of existing TT2V models for video reconstruction, we first extract the representative text constructed with our selection strategy, effectively encoding video content as the spatial component and motion as the temporal component.
 Then, the video generation model is utilized to reconstruct the corresponding video from text inputs.
-The rationale behind this strategy lies in that the compact yet effective text representations from the encoder encapsulate semantic details to enable high-quality semantic reconstruction for the decoder.
+The rationale behind this strategy lies that compact yet effective text representations from the encoder encapsulate semantic details to enable high-quality semantic reconstruction for the decoder.
 Different from TT2V, the IT2V mode is designed to enhance perceptual reconstruction, since images provide richer visual contexts compared to text, benefiting perceptual consistency.
-This is achieved by inputting similar text representations with TT2V mode and extra selected keyframes from the encoder to the decoder for better perceptual video reconstruction.
-To further ensure perceptual smoothness across consecutive frames, an efficient adaptation tuning in a frame-interpolation manner via Low-Rank Adaption (LoRA) tuning is tailored to fully exploit the semantic cues and visual contexts from both input texts and keyframes to facilitate high-quality perceptual consistency for video reconstruction.
+This is achieved by inputting similar text representations with the TT2V mode and extra selected keyframes from the encoder to the decoder for better perceptual video reconstruction.
+To further improve perceptual smoothness across consecutive frames, an efficient adaptation tuning in a frame-interpolation manner via Low-Rank Adaption (LoRA) tuning is tailored to fully exploit the semantic cues and visual contexts from both input texts and keyframes to facilitate high-quality perceptual consistency for video reconstruction.
 This comprehensive paradigm adeptly accommodates diverse modality representations within video coding, by tapping into foundational MLLMs and video generation models, which sheds light on future video coding works.
 The contributions of our work are as follows:
-  • We introduce a unified paradigm for CMVC. To the best of our knowledge, this is the first work to make utility of foundational MLLMs and video generation models for video coding.
-  • We elaborate multiple encoding-decoding modes to achieve good trade-off video reconstruction quality depending on specified decoding requirements, including TT2V mode to ensure high-quality semantic information and IT2V mode to achieve superb perceptual consistency.
-  • Extensive experiments demonstrate that our proposed CMVC pipeline obtains competitive video reconstructions on HEVC Class B, C, D, E, UVG and MCL-JCV benchmarks, while also maintaining high compression rates.
+  • To the best of our knowledge, the proposed unified paradigm for CMVC is the first to leverage foundational MLLMs and video generation models for video coding.
+  • We elaborate multiple encoding-decoding modes to achieve good trade-off video reconstruction quality for specified decoding requirements, including TT2V mode to ensure high-quality semantic information and IT2V mode to achieve superb perceptual consistency.
+  • Extensive experiments demonstrate that the proposed CMVC pipeline obtains competitive video reconstructions on HEVC Class B, C, D, E, UVG and MCL-JCV benchmarks while maintaining high compression ratios.
 Related works
@@
 2023; Yin et al.
-2023), and image-text-to-video (IT2V) (Zhang et al.
+2023), and IT2V (Zhang et al.
 2024a), among others.
@@
 2024a) adds visual descriptions for images by adopting latent interpolation adaptation training to produce smoothing transformation.
-Building on this concept, we can also enhance video generation by incorporating temporal descriptions.
+Building upon this concept, we can also enhance video generation by incorporating temporal descriptions.
 Cross-Modality Compression
@@
 2021) integrates the I2T translation model with the T2I (Text-to-Image) generation model.
-Building on this foundation, SCMC (Zhang et al.
+Based on this foundation, SCMC (Zhang et al.
 2023) introduces a scalable cross-modality compression paradigm that hierarchically represents images across different modalities, thereby enhancing both semantic and signal-level fidelity.
@@
 These models demonstrate that the integration of I2T and T2I methodologies has outperformed the most advanced visual signal codecs.
-Despite these advancements, there remains limited research focused on cross-modality video coding
-Figure 2: Visualization of the results from the keyframe selection strategy. The red points indicate the index of the keyframes. The first and last frames are set as keyframes by default.
+Despite these advancements, there remains limited research focused on cross-modality video coding.
 CMVC Scheme
 Overview
-We propose a unified paradigm for cross-modality video compression (CMVC), as illustrated in Fig.
+We propose a CMVC paradigm for efficient video compression with high semantic and perceptual quality, especially at low bitrates, as illustrated in Fig.
 1.
-First, the video is segmented into multiple clips through the keyframe selection strategy, and each clip is disentangled into the contents of keyframes and motion components.
-Subsequently, MLLMs are employed to create multimodal representations for these components.
-For instance, keyframes and motion can be represented using V2T models to convert them into textual formats.
-These representations of keyframes and motion are then compressed and transmitted to the decoder.
-We design two distinct modes based on the modalities of keyframes and motion: TT2V mode and IT2V mode.
-In TT2V mode, keyframes and motion are both represented as text, enabling the generation of videos with high semantic consistency at ULB.
-In IT2V mode, keyframes are represented as images, and motion is described using text, facilitating video generation with high perceptual consistency.
-Overall, this paradigm offers flexibility to integrate various SOTA models and mode conversions while maintaining semantic and perceptual quality at high compression ratios.
+Given a video $V\in v_{i}$, where $v_{i}$ denotes video frames and $i\in\{1,\dots,N\}$, which consists of spatial (keyframe) and temporal (motion) components, the goal is to compress these components into compact multimodal representations.
+We leverage MLLMs, specifically V2T models, to map both keyframes and motion into textual representations, which are then encoded using specialized encoders.
+These multimodal representations are then compressed using dedicated encoders, yielding compressed representations of keyframe and motion.
+The video is reconstructed by a decoder operating in one of two modes: TT2V, which prioritizes semantic consistency, and IT2V, which focuses on perceptual quality.
+This approach enables high compression ratios while preserving both semantic information and perceptual quality.
 CMVC Encoder
@@
 We calculate the cosine similarity distance between the first frame and subsequent frames as follows:
-$\centering\mathcal{D}_{v}=\frac{v_{k}\cdot v_{k+i}}{\left\|v_{k}\right\|\cdot\left\|v_{k+i}\right\|}=\frac{\sum_{j=1}^{m}v_{k,j}\cdot v_{k+i,j}}{\sqrt{\sum_{j=1}^{m}\left(v_{k,j}\right)^{2}\cdot\sum_{j=1}^{m}\left(v_{k+i,j}\right)^{2}}},\@add@centering$ (1)
-where $v_{k+i}$ is the feature vectors extracted from the subsequent frames.
-$m$ is the number of components of vectors $v_{k}$ and $v_{k+i}$.
+$\centering\mathcal{D}_{c}=\frac{c_{k}\cdot c_{k+i}}{\left\|c_{k}\right\|\cdot\left\|c_{k+i}\right\|}=\frac{\sum_{j=1}^{m}c_{k,j}\cdot c_{k+i,j}}{\sqrt{\sum_{j=1}^{m}\left(c_{k,j}\right)^{2}\cdot\sum_{j=1}^{m}\left(c_{k+i,j}\right)^{2}}},\@add@centering$ (1)
+where $c_{k+i}$ is the feature vectors extracted from the subsequent frames.
+$m$ is the number of components of vectors $c_{k}$ and $c_{k+i}$.
 Within a uniform interval, we select the frame with the smallest similarity to the previous keyframes to form a set of keyframes that better showcase significant motion.
-Then, the $v_{k}$ is replaced by the next keyframe features, acting like a dynamic mechanism.
+Then, the $c_{k}$ is replaced by the next keyframe features, acting like a dynamic mechanism.
 This iterative process is repeated for subsequent clips, systematically identifying representative keyframes.
-We display the result of the keyframe selection strategy in Fig.
-2.
 Multimodality representation.
@@
 The keyframes are denoted as $K$ = {$k_{1}$, $k_{2}$, …, $k_{n}$}, where each $k_{i}$ is a keyframe.
-The motion information $m_{j}$ between these keyframes is represented by $M$ = {$m_{1}$, $m_{2}$, …, $m_{n-1}$}.
+The motion information $m_{j}$ between two consecutive keyframes is represented by $M$ = {$m_{1}$, $m_{2}$, …, $m_{n-1}$}.
 Keyframes and motion are transformed into multimodality representations as follows: $T_{k,i}=f(k_{i})$ and $T_{m,j}=g(m_{j})$.
@@
 As illustrated in Fig.
-1, keyframes can be transformed into textual and visual image representations, while motion can be represented by textual description and optical flow.
+1, keyframes and motion can be transformed into textual and visual representations.
 Thus, the total bitrate is given by:
@@
 CMVC Decoder
-In the decoder, we utilize decoded keyframe $\hat{K}$ and motion $\hat{M}$ to achieve video generation, as follows:
+In the decoder, we utilize the decoded keyframe $\hat{K}$ and motion $\hat{M}$ to achieve video generation, as follows:
 $\hat{V}=\mathcal{G}(\hat{K},\hat{M}),$ (3)
 where $\mathcal{G}(*)$ is a video generation model and $\hat{V}$ is the reconstructed video.
-According to different modality representations for keyframe and motion, we design two modes, including TT2V mode and IT2V mode.
-In the TT2V mode, we employ SOTA video generation models, such as Open-Sora (Lab and etc.
-2024), VideoCrafter2 (Chen et al.
-2024a), AnimateDiff (Guo et al.
-2024) and etc., to generate videos based on the decoded descriptions of keyframes and motion.
-By leveraging advancements in these models, we can employ a range of techniques that lead to optimal semantic reconstruction.
-Our findings indicate that more detailed descriptions correspond to higher bitrates, resulting in better semantic quality of the generated videos.
-In the IT2V mode, we integrate keyframe images and motion descriptions to enhance the perceptual quality of the generative video.
-In addition to adopting existing IT2V models, we propose an efficient generative model utilizing LoRA tuning to ensure higher-quality perceptual consistency at ELB.
+Based on different modality representations for keyframe and motion, we designed two modes, including the TT2V mode and the IT2V mode for ULB and ELB coding, respectively.
+In the TT2V mode, we utilize state-of-the-art (SoTA) video generation models to generate videos from decoded keyframes and motion descriptions.
+Leveraging advancements in these models, we optimize semantic reconstruction, with our results showing that more detailed descriptions yield higher bitrates and improved semantic quality.
+In the IT2V mode, keyframe images and motion descriptions are integrated to enhance perceptual quality.
+In addition to employing existing IT2V models, we propose a generative model utilizing LoRA tuning to ensure superior perceptual consistency at ELB.
 The IT2V generative model.
@@
 Two LoRAs are trained to fit the two keyframe images ($I_{0}$ and $I_{1}$), respectively.
-To generate $w$-th frame between $I_{0}$ and $I_{1}$, we interpolate $I^{{}^{\prime}}_{w}$ and the LoRA parameters according to the weights $w_{i}$ and $w_{l}$, where $w_{i}$ and $w_{l}$ are optimized according to the $w$-th frame.
+To generate $w$-th frame between $I_{0}$ and $I_{1}$, we interpolate $I^{{}^{\prime}}_{w}$ and the LoRA parameters according to the weights $w_{i}$ and $w_{l}$.
 The IT2V mode is designed to obtain a reconstructed video according to keyframe images and the text of motion.
 Thus, we propose a IT2V generative model, which generates a video clip according to two keyframe images ($I_{0}$ and $I_{1}$) and the description of the motion of this video clip.
-Specifically, we adopt a stable diffusion (SD) model with LoRA, which fine-tunes the model parameters $\theta$ by training a low-rank residual component $\triangle\theta$.
+Specifically, we adopt a stable diffusion model (SD) with LoRA, which fine-tunes the model parameters $\theta$ by training a low-rank residual component $\triangle\theta$.
 This residual can be decomposed into products of low-rank matrices.
@@
 The proposed IT2V model is shown in Fig.
-3.
+2.
 We first train two LoRAs ($\Delta\theta_{0}$ and $\Delta\theta_{1}$) on the SD UNet $\epsilon_{\theta}$ for each of the two images $I_{0}$ and $I_{1}$.
@@
 Figure 3: Left: Comparison results of combination of different V2T models (VideoLLaVA and VideoLLaMA) and TT2V models (VideoCrafter1, VideoCrafter2, ModelScope, OpenSora and AnimateDiff).
-Right: Visual quality comparison of the TT2V mode and VVC.
+Right: Visual quality comparison of the TT2V mode and VTM.
 At ULB, our proposed TT2V mode successfully preserves the semantic quality of the videos.
-In contrast, VVC brings significant blocking artifacts, which impedes the effective conveyance of semantic information in videos.
+In contrast, VTM brings significant blocking artifacts, which impedes the effective conveyance of semantic information in videos.
 Figure 4: Visual quality comparison. The values represent the BPP (1e-2) and the DISTS value. A lower DISTS value indicates better perceptual quality.
@@
 Figure 5: The R-D performance comparison in the IT2V mode. The comparisons are performed on the Class B, Class C, Class D, Class E, UVG, and MCL-JCV, respectively.
-Table 1: BD-Rate (%) comparison of different video generation models across various datasets in terms of DISTS.
+Table 1: BD-Rate (%) comparison of different video generation models across various datasets in terms of DISTS. The anchor is VTM with QP={52, 50, 47, 45}.
 RIFE AMT DiffMorpher Ours
-Class B 6.00 47.90 -46.47 -65.71
-Class C 21.85 14.87 -$3$31.15
-Class D -17.82 -29.88 -51.40 -54.45
-Class E -24.61 -79.22 38.03 -90.40
-UVG 23.87 42.59 13.62 -24.21
-MCL-JCV 87.37 164.10 24.60 -19.73
+Class B 4.32 53.97 -45.94 -59.12
+Class C 42.58 50.57 -31.06 -17.27
+Class D -4.89 -29.88 -49.06 -52.16
+UVG 24.08 15.00 11.55 -21.18
+MCL-JCV 83.08 151.20 22.16 -7.25
 Comparison methods.
 There are numerous SOTA foundation models available for video understanding.
-We choose two prominent models, namely VideoLLaVA and VideoLLaMA, to extract semantic information from videos.
+We choose two prominent models, namely VideoLLaVA (Lin et al.
+2023) and VideoLLaMA (Zhang, Li, and Bing 2023), to extract semantic information from videos.
 This process aligns with the V2T stage depicted in Fig.
@@
 In the TT2V mode, numerous video generation models are available.
-In this context, we employ advanced video generation models, including Open-Sora, VideoCrafter1, VideoCrafter2, and AnimateDiff, for the purpose of generating videos based on textual input.
-In addition, we compare with the video codec VVC at the extremely low bitrate with QP=63.
-In the IT2V mode, we conduct a comparative analysis of existing traditional video codecs, such as x264, x265, and VVC (Bross et al.
+In this context, we employ advanced video generation models, including Open-Sora (Lab and etc.
+2024), VideoCrafter1 (Chen et al.
+2023), VideoCrafter2 (Chen et al.
+2024a), and AnimateDiff (Guo et al.
+2024), for the purpose of generating videos based on textual input.
+In addition, we compare with the video codec VTM at the extremely low bitrate with QP=63.
+In the IT2V mode, we conduct a comparative analysis of existing traditional video codecs, such as x264, x265, and VTM (Bross et al.
 2021).
@@
 Furthermore, it should be emphasized that the bit consumption associated with motion text has not been calculated.
-Notably, the anchor of BD-Rate is VVC under Low Delay P configuration with QPs={63, 52, 47, 42}.
 Figure 6: Visual quality comparison in a video. The numbers displayed beneath the images correspond to the frame index.
@@
 Comparison in the TT2V mode.
-We conduct a comparative analysis of two V2T models, VideoLLaVA and VideoLLaMA, both of which are state-of-the-art (SOTA) MLLMs.
+We conduct a comparative analysis of two V2T models, VideoLLaVA and VideoLLaMA, both of which are the SOTA MLLMs.
 Subsequently, we extend our comparison to include five video generation models: VideoCrafter1, VideoCrafter2, ModelScope, OpenSora, and AnimateDiff.
-Furthermore, we compare these models against the traditional video codec VVC at QP=63, which results in a higher bitrate than our proposed scheme.
+Furthermore, we compare these models against the traditional video codec VTM at QP=63, which results in a higher bitrate than our proposed scheme.
 Our assessment focuses on five aspects: subject consistency, background consistency, temporal flickering, motion smoothness, and frame quality.
 The results, illustrated in Fig.
-4, indicate that the TT2V generation models outperformed VVC, showcasing better frame-wise quality and consistency in both background and subject representation.
+3, indicate that the TT2V generation models outperformed VTM, showcasing better frame-wise quality and consistency in both background and subject representation.
 These results reflect the average performance across all testing datasets, and detailed comparison results can be found in the supplementary material.
 The visual quality comparisons illustrated in Fig.
-4 indicate that VVC displays considerable blocking artifacts, which severely hinder its ability to convey semantic information.
+3 indicate that VTM displays considerable blocking artifacts, which severely hinder its ability to convey semantic information.
 Table 2: Ablation studies on the keyframe selection strategy.
@@
 Comparison in the IT2V mode.
-We compare our model with traditional codecs (x264, x265, and VVC) as well as deep video codecs (DCVC and DCVC-DC).
+We compare our model with traditional codecs (x264, x265, and VTM) as well as deep video codecs (DCVC and DCVC-DC).
 As presented in Fig.
-6, we use DISTS to evaluate perceptual quality.
+5, we use DISTS to evaluate perceptual quality.
 Additional comparisons with other evaluation metrics, such as LPIPS, FID, and PSNR, are provided in the supplementary material.
@@
 4 and Fig.
-7.
+6.
 Our proposed model exhibits superior perceptual quality in both spatial and temporal dimensions.
@@
 2018), NIC (Chen et al.
-2021), and NTVP.
-Among these, NTVP stands out by demonstrating superior reconstructed quality while maintaining a lower coding rate.
+2021), and NVTC (Feng et al.
+2023).
+Among these, NVTC stands out by demonstrating superior reconstructed quality while maintaining a lower coding rate.
 Our model requires updating $w_{i}$ and $w_{l}$ based on the input, such that we further evaluate the effectiveness of updating strategies, as illustrated in Table 3.
@@
 100 -25.207
-Conclusion and Discussion
-We propose a CMVC paradigm that represents a promising advancement in video coding technology.
-This framework effectively tackles the challenges of preserving semantic integrity and perceptual consistency at ULB and ELB.
-By leveraging MLLMs and cross-modality representation techniques, the proposed CMVC framework disentangles videos into content and motion components, transforming them into different modalities for efficient compression and reconstruction.
-Through the TT2V and IT2V modes, CMVC achieves a balance between semantic information and perceptual quality, offering a comprehensive solution at high compression ratios.
+Discussion
+Application.
+The proposed method enables efficient transmission at low bitrates while preserving semantic content, making it ideal for bandwidth-limited or emergency alert scenarios.
+When bandwidth allows, keyframe data can be transmitted with textual descriptions, allowing the decoder to reconstruct an approximate video.
+This hybrid approach balances visual quality with data efficiency, suitable for situations where full video streaming is infeasible, such as disaster response and remote surveillance.
+Although still in the research phase, the method shows strong potential for real-world applications.
+With advances in computational resources and model optimization techniques like pruning and quantization, it is expected to become a practical solution for emergency communications and other bandwidth-constrained environments.
+Higher bitrate.
 CMVC includes the TT2V mode and IT2V mode, but more modes can be further explored.
@@
 This approach aims to achieve superior performance across all bitrates and dimensions when compared to traditional codecs.
+Conclusion
+We propose a CMVC paradigm that represents a promising advancement in video coding technology.
+This framework effectively tackles the challenges of preserving semantic integrity and perceptual consistency at ULB and ELB.
+By leveraging MLLMs and cross-modality representation techniques, the proposed CMVC framework disentangles videos into content and motion components, transforming them into different modalities for efficient compression and reconstruction.
+Through the TT2V and IT2V modes, CMVC achieves a balance between semantic information and perceptual quality, offering a comprehensive solution at high compression ratios.